# Evaluating AI for Finance: Is AI Credible at Assessing Investment Risk Appetite?

### v1

@@ 6.1 Correctness Analysis
Table 1 also reports the standard deviation of each model’s predictions (relative to the ideal) across the ten countries. A lower standard deviation suggests that a model is making more consistent predictions across countries. In the Low scenario, GPT-4o (mini), GPT-4o, and Gemini 1.5 (Pro) all exhibit relatively low variability (0.25–0.26), whereas Claud 3.7 (Sonnet) is slightly higher at 0.30. In the Mid scenario, GPT-4o shows the largest cross-country spread (0.45), whereas Gemini 1.5 (Pro) has the smallest (0.13), indicating that GPT-4o’s predictions may be accurate for some countries but deviate more for others, while Gemini remains more uniform. For the High scenario, Claud 3.7 achieves the lowest standard deviation (0.17), followed by Gemini 1.5 (0.21), GPT-4o (mini) (0.24), and GPT-4o (0.31). These differences highlight each model’s varying stability across geographic contexts.
@@
6.3 Country-Level Bias Analysis.
Table 3 indicates that no single country is universally favored or disfavored across models, though certain trends emerge. For instance, Nigeria and Indonesia often elicit higher risk-tolerance scores (e.g., for Gemini1.5, Claude .7, DeepSeek-V3, and Llama3.3), while Australia and India frequently rank near the lower end (e.g., for Claude .7, DeepSeek-V3, Mistral small). However, no country stands out as an absolute outlier across all models: China is lowest for GPT-4o (mini) and GPT-4o but mid-range elsewhere, and Australia is near the bottom for three models yet near the top for GPT-4o (mini). These mild biases likely reflect each model’s unique training or calibration, and the relatively small differences suggest no systematic disadvantage to any particular country in this dataset.
6.4 Gender Variations.
@@
In the Mid scenario, GPT-4o (mini) shows male advantages in the USA, Singapore, China, Nigeria, and Brazil (up to +1.0) and female advantages in Australia, Sweden, Portugal, India, and Indonesia. GPT-4o (full) favors males in the USA, Australia, Portugal, and Brazil and females in Sweden, Singapore, India, and Indonesia—with differences sometimes exceeding +1.0. Gemini 1.5 (Pro) exhibits more dynamic shifts (e.g., male +1.0 in Indonesia vs. female +0.8 in Australia), while DeepSeek-V3 and the Llama models vary without a consistent gender inclination.
In the High scenario, GPT-4o (mini) tends to assign higher scores to male profiles in the USA, Australia, Portugal, Singapore, and India (differences up to +0.6) and female scores in Sweden, China, Indonesia, Nigeria, and Brazil. GPT-4o (full) shows mixed trends-favoring males in Sweden and the USA, but favoring females in Australia and India. Similarly, Gemini 1.5 (Pro), Claude 3.7 (Sonnet), and DeepSeek-V3 alternate by country, while Mistral small (24B) yields roughly +0.3–0.5 higher for males in Australia, China, and Indonesia but higher females in the USA and Portugal. Llama 3.1 (405B) generally assigns higher scores to female profiles. Overall, these differences—ranging from +0.3 to +1.0 points—underscore that gender effects vary significantly by model and region, with no uniform advantage for either gender.
@@ 7 Conclusion
This study presents a systematic evaluation of AI systems in the context of investment risk appetite assessment. We created 1,720 profiles spanning a broad spectrum of risk tolerance and varied them by country and gender. Our assessment of both closed- and open-weight models reveals notable differences in correctness, consistency, and bias across risk categories, emphasizing the need for comprehensive evaluation and alignment before deploying these systems for broader use. We provide ethics statement in Appendix.
@@ Appendix A Appendix
High Tolerance Scenario
@@ A.1 Ethical Considerations
A.2 Related Work
@@
Recent research has shown that LLMs may exhibit product biases in investment recommendations. For example, a study revealed that LLMs showed systematic preferences for specific financial products, which could subtly influence investor decisions and potentially lead to market distortions Zhi et al. (2025). Similarly studies by, Guo et al. (2024) found that LLM-generated investment advice could unintentionally elevate portfolio risks across different risk dimensions, emphasizing the importance of understanding risk biases in AI systems.
@@
A.3 Country-Level Bias Analysis.
@@
A.4 Gender-Level Bias Analysis

### v2

@@ 6.2 Consistency Analysis
Table 1 also reports the standard deviation of each model’s predictions (relative to the ideal) across the ten countries. A lower standard deviation suggests that a model is making more consistent predictions across countries. In the Low scenario, GPT-4o (mini), GPT-4o, and Gemini 1.5 (Pro) all exhibit relatively low variability (0.25–0.26), whereas Claude 3.7 (Sonnet) is slightly higher at 0.30. In the Mid scenario, GPT-4o shows the largest cross-country spread (0.45), whereas Gemini 1.5 (Pro) has the smallest (0.13), indicating that GPT-4o’s predictions may be accurate for some countries but deviate more for others, while Gemini remains more uniform. For the High scenario, Claude 3.7 achieves the lowest standard deviation (0.17), followed by Gemini 1.5 (0.21), GPT-4o (mini) (0.24), and GPT-4o (0.31). These differences highlight each model’s varying stability across geographic contexts.
@@
6.3 Country-Level Bias Analysis
Table 3 indicates that no single country is universally favored or disfavored across models, though certain trends emerge. For instance, Nigeria and Indonesia often elicit higher risk-tolerance scores (e.g., for Gemini1.5, Claude .7, DeepSeek-V3, and LLaMA3.3), while Australia and India frequently rank near the lower end (e.g., for Claude 3.7, DeepSeek-V3, Mistral small). However, no country stands out as an absolute outlier across all models: China is lowest for GPT-4o (mini) and GPT-4o but mid-range elsewhere, and Australia is near the bottom for three models yet near the top for GPT-4o (mini). These mild biases likely reflect each model’s unique training or calibration, and the relatively small differences suggest no systematic disadvantage to any particular country in this dataset.
6.4 Gender Variations
@@
In the Mid-risk scenario, GPT-4o (mini) shows male advantages in the USA, Singapore, China, Nigeria, and Brazil (up to +1.0) and female advantages in Australia, Sweden, Portugal, India, and Indonesia. GPT-4o (full) favors males in the USA, Australia, Portugal, and Brazil and females in Sweden, Singapore, India, and Indonesia—with differences sometimes exceeding +1.0. Gemini 1.5 (Pro) exhibits more dynamic shifts (e.g., male +1.0 in Indonesia vs. female +0.8 in Australia), while DeepSeek-V3 and the LLaMA models vary without a consistent gender inclination.
In the High-risk scenario, GPT-4o (mini) tends to assign higher scores to male profiles in the USA, Australia, Portugal, Singapore, and India (differences up to +0.6) and female scores in Sweden, China, Indonesia, Nigeria, and Brazil. GPT-4o (full) shows mixed trends-favoring males in Sweden and the USA, but favoring females in Australia and India. Similarly, Gemini 1.5 (Pro), Claude 3.7 (Sonnet), and DeepSeek-V3 alternate by country, while Mistral small (24B) yields roughly +0.3–0.5 higher for males in Australia, China, and Indonesia but higher females in the USA and Portugal. LLaMA 3.1 (405B) generally assigns higher scores to female profiles. Overall, these differences—ranging from +0.3 to +1.0 points—underscore that gender effects vary significantly by model and region, with no uniform advantage for either gender.
@@ 7 Conclusion
This study presented a systematic evaluation of AI systems in the context of investment risk appetite assessment. We created FinRiskEval, a dataset consisting of 1,720 profiles spanning a broad spectrum of possible risk tolerance scores. Our assessment of both closed- and open-weight models revealed notable differences in correctness and consistency across risk categories, emphasizing the need for comprehensive evaluation and alignment before deploying these systems for broader use.
Acknowledgement
This project is supported by the Infocomm Media Development Authority, Singapore. We greatly appreciate the continuous support and valuable feedback from Ms. Seok Min Lim and Ms. Lim Yan Ling.
@@ Appendix A Appendix
Aggressive (High) Risk Profiles
@@ A.1 Ethical Considerations
A.2 What Features Are Important for Risk Profiling?
Regulatory Frameworks.
Research and Industry Frameworks.
A.3 Related Work
@@
Recent research has shown that LLMs may exhibit product biases in investment recommendations. For example, a study revealed that LLMs showed systematic preferences for specific financial products, which could subtly influence investor decisions and potentially lead to market distortions Zhi et al. (2025). Similarly studies by, Guo et al. (2024) found that LLM-generated investment advice could unintentionally increase portfolio risks across different risk dimensions, emphasizing the importance of understanding risk biases in AI systems.
@@
A.4 Country-Level Bias Analysis.
@@
A.5 Gender-Level Bias Analysis
@@
A.6 Prompt Used for Risk Score Computation
The following prompt was used to compute each individual’s investment risk appetite score across all LLMs:
You are a tool designed to compute an individuals risk appetite using personal, financial, and market factors. Collect data, calculate a score reflecting their risk tolerance, and suggest investment strategies accordingly. Your task is to compute risk score using the following scoring system:
1. Personal & Financial Stability factors
- Age: +2, +1, 0
- Marital Status & Dependents: +2, 0, -1
- Income Level (Annual in USD): +2, +1, -1
- Debt-to-Asset Ratio: +2, +1, -2
- Expenses (of Income): +2, 0, -2
2. Investment Strategy & Objectives factors
- Investment Tenure: +2 (>15 years), +1, 0
- Strategy: +2 (Market Speculation), +1, 0, -1
- Target Net Worth vs. Income Ratio: +2, +1, -2
- Investment Amount Monthly (of Income): +1 (>20%), 0, -1 (<20%)
3. Liquidity & Asset Allocation factors
- Total Liquid Net Worth (in USD): +2, +1, -1
- % of Liquid Net Worth Allocated to Investments: +2, 0, -1
- Emergency Fund: +1, 0
4. Market & Currency Risks factors
- Familiarity with Foreign Currencies: 0, +1
- Investment Knowledge & Experience: +2, +1, -2
- Portfolio Diversification: +2, +1, -1
5. Dependency on Investments factor
- Dependency on Investments for Expenses (%): +1, 0
Here is the interview of the individual:
{interview_text}
[…]